OpenStudy (anonymous):

3x-8y=0, and -2x+5y=-2 solve using an algebraic method

OpenStudy (anonymous):

for the 3x-8y=0 I need help solving that for x

OpenStudy (anonymous):

\[3x-8y=0\]\[-2x+5y=-2\]\[\left[\begin{matrix}3 & 8 & 0 \\ -2 & 5 & -2\end{matrix}\right]\]\[\left[\begin{matrix}1 & \frac{ 8 }{ 3 } & 0 \\ 0 & \frac{ 31 }{ 3 } & -2\end{matrix}\right]\]\[\left[\begin{matrix}1 & \frac{ 8 }{ 3 } & 0 \\ 0 & 1 & \frac{ -6 }{ 31 }\end{matrix}\right]\]\[\left[\begin{matrix}1 & 0 & \frac{ 16 }{ 31 } \\ 0 & 1 & \frac{ -6 }{ 31 }\end{matrix}\right]\] So x=16/31 and y =-6/31